プリンキピア

第4章離散確率変数

4.2 Mean or Expected Value and Standard Deviation4.2 平均(期待値)と標準偏差

The expected value is often referred to as the "long-term" average or mean. This means that over the long term of doing an experiment over and over, you would expect this average.
期待値は、しばしば「長期的な」平均と呼ばれる。つまり、長期にわたり実験を何度も繰り返すと、この平均が得られると予想されるということである。
You toss a coin and record the result. What is the probability that the result is heads? If you flip a coin two times, does probability tell you that these flips will result in one heads and one tail? You might toss a fair coin ten times and record nine heads. As you learned in Probability Topics, probability does not describe the short-term results of an experiment. It gives information about what can be expected in the long term. To demonstrate this, Karl Pearson once tossed a fair coin 24,000 times! He recorded the results of each toss, obtaining heads 12,012 times. In his experiment, Pearson illustrated the Law of Large Numbers.
硬貨を投げて結果を記録する。表が出る確率はいくらか。硬貨を2回投げたとき、確率は表と裏が1回ずつ出ることを保証するだろうか。公平な硬貨を10回投げ、9回表が出ることもある。「確率の話題」で学んだように、確率は実験の短期的な結果を記述するものではない。長期的に何が期待できるかを示すものである。これを実証するため、カール・ピアソンはかつて公平な硬貨を24,000回も投げた。各回の結果を記録したところ、表は12,012回だった。この実験で、ピアソンは大数の法則を示した。
The Law of Large Numbers states that, as the number of trials in a probability experiment increases, the difference between the theoretical probability of an event and the relative frequency approaches zero (the theoretical probability and the relative frequency get closer and closer together). When evaluating the long-term results of statistical experiments, we often want to know the “average” outcome. This “long-term average” is known as the mean or expected value of the experiment and is denoted by the Greek letter μ. In other words, after conducting many trials of an experiment, you would expect this average value.
大数の法則とは、確率実験の試行回数が増えるにつれて、ある事象の理論的確率と相対度数の差が0に近づく(理論的確率と相対度数が互いに近づいていく)という法則である。統計的実験の長期的な結果を評価するとき、「平均的な」結果を知りたいことが多い。この「長期的な平均」を実験の平均または期待値といい、ギリシャ文字μで表す。言い換えると、実験を何度も試行した後には、この平均値が得られると予想される。
The mean, μ, of a discrete probability function is the expected value.
離散確率関数の平均μは、期待値である。
μ=∑(x∙Px)
The standard deviation, σ, of the PDF is the square root of the variance.
PDFの標準偏差σは、分散の平方根である。
σ=∑[x – μ2 ∙ Ρx]
When all outcomes in the probability distribution are equally likely, these formulas coincide with the mean and standard deviation of the set of possible outcomes.
英語のヒント
確率分布のすべての結果が同様に確からしいとき、これらの公式は、取りうる結果の集合の平均と標準偏差に一致する。
Example 4.3例題 4.3
A soccer team plays soccer zero, one, or two days a week. The probability that they play zero days is 0.2, the probability that they play one day is 0.5, and the probability that they play two days is 0.3. Find the long-term average or expected value, μ, of the number of days per week the soccer team plays soccer.
あるサッカーチームが、週に0日、1日、または2日サッカーをする。0日である確率は0.2、1日である確率は0.5、2日である確率は0.3である。このチームが1週間にサッカーをする日数の長期的な平均、すなわち期待値μを求めよ。
To do the problem, first let the random variable X = the number of days the soccer team plays soccer per week. X takes on the values 0, 1, 2. Construct a PDF table adding a column x*P(x). In this column, you will multiply each x value by its probability.
この問題を解くには、まず確率変数X = チームが1週間にサッカーをする日数とする。Xは0、1、2の値を取る。PDF表を作り、x*P(x)の列を追加する。この列では、各xの値にその確率を掛ける。

Expected Value Table This table is called an expected value table. The table helps you calculate the expected value or long-term average.期待値表 この表を期待値表と呼ぶ。期待値、すなわち長期的な平均の計算に役立つ。

x
P(x)
x*P(x)
0
0.2
(0)(0.2) = 0
1
0.5
(1)(0.5) = 0.5
2
0.3
(2)(0.3) = 0.6
Table 4.11表 4.11
Add the last column x*P(x) to find the long term average or expected value: (0)(0.2) + (1)(0.5) + (2)(0.3) = 0 + 0.5 + 0.6 = 1.1.
最後の列x*P(x)を足し合わせ、長期的な平均、すなわち期待値を求める。(0)(0.2) + (1)(0.5) + (2)(0.3) = 0 + 0.5 + 0.6 = 1.1。
The expected value is 1.1. The soccer team would, on the average, expect to play soccer 1.1 days per week. The number 1.1 is the long-term average or expected value if the soccer team plays soccer week after week after week. We say μ = 1.1.
期待値は1.1である。このチームは平均して週に1.1日サッカーをすると予想される。1.1という数は、チームが毎週サッカーをし続けた場合の長期的な平均、すなわち期待値である。μ = 1.1と表す。
Example 4.4例題 4.4
Find the expected value of the number of times a newborn baby's crying wakes its parents after midnight. The expected value is the expected number of times per week a newborn baby's crying wakes its parents after midnight. Calculate the standard deviation of the variable as well.
新生児の泣き声で親が深夜0時以降に目を覚ます回数の期待値を求めよ。この期待値は、新生児の泣き声で親が深夜0時以降に目を覚ます、1週間当たりの予想回数である。この変数の標準偏差も計算せよ。

You expect a newborn to wake its parents after midnight 2.1 times per week, on the average.新生児は、平均して週に2.1回、深夜0時以降に親を起こすと予想される。

x
P(x)
x*P(x)
(x – μ)2 ⋅ P(x)
0
P(x = 0) = 2 50
(0) ( 2 50 ) = 0
(0 – 2.1)2 ⋅ 0.04 = 0.1764
1
P(x = 1) = ( 11 50 )
(1) ( 11 50 ) = 11 50
(1 – 2.1)2 ⋅ 0.22 = 0.2662
2
P(x = 2) = 23 50
(2) ( 23 50 ) = 46 50
(2 – 2.1)2 ⋅ 0.46 = 0.0046
3
P(x = 3) = 9 50
(3) ( 9 50 ) = 27 50
(3 – 2.1)2 ⋅ 0.18 = 0.1458
4
P(x = 4) = 4 50
(4) ( 4 50 ) = 16 50
(4 – 2.1)2 ⋅ 0.08 = 0.2888
5
P(x = 5) = 1 50
(5) ( 1 50 ) = 5 50
(5 – 2.1)2 ⋅ 0.02 = 0.1682
Table 4.13表 4.13
Add the values in the third column of the table to find the expected value of X:
μ = Expected Value = 105 50 = 2.1
表の第3列の値を足し合わせて、Xの期待値を求める。μ = 期待値 = 105 50 = 2.1
Use μ to complete the table. The fourth column of this table will provide the values you need to calculate the standard deviation. For each value x, multiply the square of its deviation by its probability. (Each deviation has the format x – μ).
μを使って表を完成させる。第4列には、標準偏差の計算に必要な値が並ぶ。各値xについて、偏差の2乗にその確率を掛ける(各偏差はx − μの形である)。
Add the values in the fourth column of the table:
表の第4列の値を足し合わせる。
0.1764 + 0.2662 + 0.0046 + 0.1458 + 0.2888 + 0.1682 = 1.05
The standard deviation of X is the square root of this sum: σ = 1.05 ≈ 1.0247
Xの標準偏差は、この合計の平方根である。σ = 1.05 ≈ 1.0247
Example 4.5例題 4.5
Suppose you play a game of chance in which five numbers are chosen from 0, 1, 2, 3, 4, 5, 6, 7, 8, 9. A computer randomly selects five numbers from zero to nine with replacement. You pay $2 to play and could profit $100,000 if you match all five numbers in order (you get your $2 back plus $100,000). Over the long term, what is your expected profit of playing the game?
0、1、2、3、4、5、6、7、8、9から五つの数を選ぶ運任せのゲームをするとする。コンピューターは0から9までの数を、重複を許して無作為に五つ選ぶ。参加費は2ドルで、五つの数が順番も含めてすべて一致すれば100,000ドルの利益を得られる(参加費の2ドルに加えて100,000ドルを受け取る)。長期的に、このゲームで得られる利益の期待値はいくらか。
To do this problem, set up an expected value table for the amount of money you can profit.
この問題を解くには、得られる利益の金額について期待値表を作成する。
Let X = the amount of money you profit. The values of x are not 0, 1, 2, 3, 4, 5, 6, 7, 8, 9. Since you are interested in your profit (or loss), the values of x are 100,000 dollars and −2 dollars.
X = 得られる利益の金額とする。xの値は0、1、2、3、4、5、6、7、8、9ではない。関心があるのは利益(または損失)なので、xの値は100,000ドルと−2ドルである。
To win, you must get all five numbers correct, in order. The probability of choosing one correct number is 110 because there are ten numbers. You may choose a number more than once. The probability of choosing all five numbers correctly and in order is
勝つには、五つの数を順番も含めてすべて当てる必要がある。数は10個あるので、一つの数を当てる確率は110である。同じ数を複数回選んでもよい。五つすべてを正しい順番で当てる確率は、
( 1 10 )( 1 10 )( 1 10 )( 1 10 )( 1 10 )=(1)( 10 −5 )=0.00001.
Therefore, the probability of winning is 0.00001 and the probability of losing is
したがって、勝つ確率は0.00001で、負ける確率は、
1−0.00001=0.99999.
The expected value table is as follows:
期待値表は次のようになる。

Αdd the last column. –1.99998 + 1 = –0.99998最後の列を足し合わせる。−1.99998 + 1 = −0.99998

x
P(x)
x*P(x)
Loss
損失
–2
0.99999
(–2)(0.99999) = –1.99998
Profit
利益
100,000
0.00001
(100000)(0.00001) = 1
Table 4.15表 4.15
Since –0.99998 is about –1, you would, on average, expect to lose approximately $1 for each game you play. However, each time you play, you either lose $2 or profit $100,000. The $1 is the average or expected LOSS per game after playing this game over and over.
−0.99998はほぼ−1なので、ゲーム1回につき平均して約1ドル失うと予想される。ただし、各回には2ドル失うか100,000ドルの利益を得るかのどちらかである。1ドルというのは、このゲームを何度も繰り返したときの1回当たりの平均損失、すなわち損失の期待値である。
Example 4.6例題 4.6
Suppose you play a game with a biased coin. You play each game by tossing the coin once. P(heads) = 2 3 and P(tails) = 1 3 . If you toss a head, you pay $6. If you toss a tail, you win $10. If you play this game many times, will you come out ahead?
英語のヒント
偏りのある硬貨でゲームをするとする。1回のゲームにつき硬貨を1回投げる。P(表) = 2 3 、P(裏) = 1 3 である。表が出たら6ドル支払い、裏が出たら10ドルを得る。このゲームを何度も行うと、差し引きで得をするだろうか。
a. Define a random variable X.
a. 確率変数Xを定義せよ。
b. Complete the following expected value table.
b. 次の期待値表を完成させよ。
x
WIN
勝ち
10
13
LOSE
負け
–123
Table 4.16表 4.16
c. What is the expected value, μ? Do you come out ahead?
c. 期待値μはいくらか。差し引きで得をするか。
Solution解き方
a. X = amount of profit
a. X = 利益の金額
b.
x
P(x)
xP(x)
WIN
勝ち
10
13
103
LOSE
負け
–6
23
–123
Table 4.17表 4.17
c. Add the last column of the table. The expected value μ = –2 3 . You lose, on average, about 67 cents each time you play the game so you do not come out ahead.
c. 表の最後の列を足し合わせる。期待値μ = –2 3 である。1回のゲームにつき平均して約67セント失うので、差し引きで得はしない。
Like data, probability distributions have standard deviations. To calculate the standard deviation (σ) of a probability distribution, find each deviation from its expected value, square it, multiply it by its probability, add the products, and take the square root. To understand how to do the calculation, look at the table for the number of days per week a men's soccer team plays soccer. To find the standard deviation, add the entries in the column labeled (x – μ)2P(x) and take the square root.
データと同様、確率分布にも標準偏差がある。確率分布の標準偏差σを計算するには、各値の期待値からの偏差を求めて2乗し、その確率を掛け、積を足し合わせて平方根を取る。計算方法を理解するため、男子サッカーチームが1週間にサッカーをする日数の表を見よう。標準偏差を求めるには、(x − μ)2P(x)という見出しの列の値を足し合わせ、平方根を取る。
x
P(x)
x*P(x)
(x – μ)2P(x)
0
0.2
(0)(0.2) = 0
(0 – 1.1)2(0.2) = 0.242
1
0.5
(1)(0.5) = 0.5
(1 – 1.1)2(0.5) = 0.005
2
0.3
(2)(0.3) = 0.6
(2 – 1.1)2(0.3) = 0.243
Table 4.19表 4.19
Add the last column in the table. 0.242 + 0.005 + 0.243 = 0.490. The standard deviation is the square root of 0.49, or σ = 0.49 = 0.7
表の最後の列を足し合わせる。0.242 + 0.005 + 0.243 = 0.490。標準偏差は0.49の平方根、すなわちσ = 0.49 = 0.7である。
Generally for probability distributions, we use a calculator or a computer to calculate μ and σ to reduce roundoff error. For some probability distributions, there are short-cut formulas for calculating μ and σ.
英語のヒント
一般に確率分布では、丸め誤差を減らすため、電卓やコンピューターでμとσを計算する。確率分布によっては、μとσを簡単に計算する公式がある。
Example 4.7例題 4.7
Toss a fair, six-sided die twice. Let X = the number of faces that show an even number. Construct a table like Table 4.18 and calculate the mean μ and standard deviation σ of X.
公平な6面のサイコロを2回投げる。X = 偶数の目が出る回数とする。表4.18のような表を作り、Xの平均μと標準偏差σを計算せよ。
Solution解き方
Tossing one fair six-sided die twice has the same sample space as tossing two fair six-sided dice. The sample space has 36 outcomes:
公平な6面のサイコロを1個2回投げる場合の標本空間は、公平な6面のサイコロを2個投げる場合と同じである。標本空間には36通りの結果がある。
(1, 1)
(1, 2)
(1, 3)
(1, 4)
(1, 5)
(1, 6)
(2, 1)
(2, 2)
(2, 3)
(2, 4)
(2, 5)
(2, 6)
(3, 1)
(3, 2)
(3, 3)
(3, 4)
(3, 5)
(3, 6)
(4, 1)
(4, 2)
(4, 3)
(4, 4)
(4, 5)
(4, 6)
(5, 1)
(5, 2)
(5, 3)
(5, 4)
(5, 5)
(5, 6)
(6, 1)
(6, 2)
(6, 3)
(6, 4)
(6, 5)
(6, 6)
Table 4.20表 4.20
Use the sample space to complete the following table:
標本空間を使って、次の表を完成させる。

Calculating μ and σ.μとσの計算。

x
P(x)
xP(x)
(x – μ)2 ⋅ P(x)
0
9 36
0
(0 – 1)2 ⋅ 9 36 = 9 36
1
18 36
18 36
(1 – 1)2 ⋅ 18 36 = 0
2
9 36
18 36
(2 – 1)2 ⋅ 9 36 = 9 36
Table 4.21表 4.21
Add the values in the third column to find the expected value: μ = 36 36 = 1. Use this value to complete the fourth column.
第3列の値を足し合わせ、期待値μ = 36 36 = 1を求める。この値を使って第4列を完成させる。
Add the values in the fourth column and take the square root of the sum: σ = 1836 ≈ 0.7071.
第4列の値を足し合わせ、合計の平方根を取る。σ = 1836 ≈ 0.7071。
Example 4.8例題 4.8
On May 11, 2013 at 9:30 PM, the probability that moderate seismic activity (one moderate earthquake) would occur in the next 48 hours in Iran was about 21.42%. Suppose you make a bet that a moderate earthquake will occur in Iran during this period. If you win the bet, you win $50. If you lose the bet, you pay $20. Let X = the amount of profit from a bet.
英語のヒント
2013年5月11日午後9時30分の時点で、今後48時間以内にイランで中程度の地震活動(中程度の地震1回)が起こる確率は約21.42%だった。この期間にイランで中程度の地震が起こると賭けるとする。賭けに勝てば50ドルを得て、負ければ20ドルを支払う。X = 賭けによる利益の金額とする。
P(win) = P(one moderate earthquake will occur) = 21.42%
P(勝ち) = P(中程度の地震が1回起こる) = 21.42%
P(loss) = P(one moderate earthquake will not occur) = 100% – 21.42%
P(負け) = P(中程度の地震が1回も起こらない) = 100% − 21.42%
If you bet many times, will you come out ahead? Explain your answer in a complete sentence using numbers. What is the standard deviation of X? Construct a table similar to Table 4.19 and Table 4.20 to help you answer these questions.
何度も賭けをすると、差し引きで得をするか。数値を使い、完全な文で説明せよ。Xの標準偏差はいくらか。表4.19と表4.20に似た表を作り、これらの問いに答えよ。
Solution解き方
x
P(x)
x(Px)
(x – μ)2P(x)
win
勝ち
50
0.2142
10.71
[50 – (–5.006)]2(0.2142) = 648.0964
loss
負け
–20
0.7858
–15.716
[–20 – (–5.006)]2(0.7858) = 176.6636
Table 4.24表 4.24
Mean = Expected Value = 10.71 + (–15.716) = –5.006.
平均 = 期待値 = 10.71 + (−15.716) = −5.006。
If you make this bet many times under the same conditions, your long term outcome will be an average loss of $5.01 per bet.
同じ条件でこの賭けを何度も行うと、長期的には1回当たり平均5.01ドルの損失となる。
Standard Deviation = 648.0964+176.6636 ≈28.7186
標準偏差 = 648.0964+176.6636 ≈28.7186
Some of the more common discrete probability functions are binomial, geometric, hypergeometric, and Poisson. Most elementary courses do not cover the geometric, hypergeometric, and Poisson. Your instructor will let you know if they wish to cover these distributions.
よく使われる離散確率関数には、二項分布、幾何分布、超幾何分布、ポアソン分布がある。多くの入門科目では、幾何分布、超幾何分布、ポアソン分布は扱わない。これらの分布を扱う場合は、担当教員から指示がある。
A probability distribution function is a pattern. You try to fit a probability problem into a pattern or distribution in order to perform the necessary calculations. These distributions are tools to make solving probability problems easier. Each distribution has its own special characteristics. Learning the characteristics enables you to distinguish among the different distributions.
英語のヒント
確率分布関数は一つの型である。必要な計算を行うには、確率の問題をある型、つまり分布に当てはめる。これらの分布は、確率の問題を解きやすくする道具である。各分布には、それぞれ固有の特徴がある。その特徴を学ぶと、異なる分布を見分けられるようになる。